You are 21-years-old and working in the marketing department of a medium-sized corporation. You are earning an annual salary of $40,000 paid every two weeks. Your employer provides a 401 (k) plan, and matches employee contributions by 50% up to a maximum 3% of your annual salary. You are in a 25% marginal tax rate.

1)Assume that you contribute 6% of your gross income to your 401 (k) account. How much will you contribute annually and per pay period to your retirement account?

a) $ 2400 and $92

b) $ 2400 and $400

c) $ 2400 and $200

2) Given your employer's contribution matching program, a total of ______ will be deposited each pay period. This will lead to a tax savings of ______ in income taxes per year.

3) All other things being equal, what is the best type of investment income? a) Tax-sheltered income b) Tax-exempt income c) Tax-deferred income d) Taxable income
Answer #1

Let the person be X

Mr X is 21 years and earning $ 40000 he is paid by company every 2  weeks so his every week income

a year has 52 weeks

So Mr x per week income =$ 40000/52

= 769.23

Mr x paid after every 2 weeks = 769.23*2

= 1538.46

(1) contribution of Mr x for 401(k) plan =6% annual

so $40000 of 6 %

= $2400

So pay per day  contribution = 2400/52*2

=92.30

=92

So Point A is correct which is 2400& 92

(2) employer deposit 50% of employee contribution but maximum 3% of Mr x salary

maximum contribution of employer = 40000*3%

= 1200

And 50% of employee contribution = $2400*50%

= $1200

So employer contribute $ 1200 annual in 401(k) plan of Mr x

And per pay period he will contribute = 1200/52*2

=46.15 i.e. $46

As Mr x contribute $ 2400 yearly in 401(k) plan so tax saving per year = 2400*25%

= $600

( 25% is given marginal rate )

answer

So pay per period contribution of employer = $46 and tax saving per year = $600

(3)(a) Tax sheltered Income is a income which is taxable in normal case but due to its non cash deduction example depreciation it is not taxed .

it can be explain by example like we receive rent from rental property and and property is asset so it is depreciated so rental income is tax sheltered by depreciation .

Tax sheltered income cannot be best investment palm as tax sheltered income are mainly available to self employed or corporation not to salaried person

(b) tax exempt income means income on which tax is not computed it can be either case of individual or corporation , some incomes arrived by certain sources which are mentioned in federal tax laws are exempt .example some coupons issued by municipal bonds are exempt.

it cannot be best investment plan as very limit sources of income are exempt

(c)Tax deferred income are income which a person earns but doesn't receive it because it is invested some where and it will not be paid till the time it is matured , like money invested in 401(k) plan is not taxed and interest earn on that is not taxed and money invested in it provide tax savings

tax deferred is best investment plan because interest earned on tax deferred plan are non taxable and income invested every year saves lots of tax

(d) Taxable income is income which is earned Any a person through various sources and is liable to give federal income tax on it for example wages ,interest revenue

Taxable income cannot be the best investment plan as income is fully taxable

so (c) tax deferred income is best plan.